# A Survey on Speech Large Language Models for Understanding

### v5

@@ III-A1 Modular Architecture
Early efforts in speech processing largely followed a modular architecture. For example, in ASR, the pipeline typically involved a separately trained acoustic model (AM), language model (LM), and decoder, integrated via weighted finite-state transducers (WFSTs). These components were optimized independently and lacked task adaptability. Typical implementations include Gaussian Mixture Model - Hidden Markov Model (GMM-HMM) frameworks and later Deep Neural Network (DNN)-HMM hybrids [6, 42], which formed the foundation of many large-scale ASR toolkits such as Kaldi [43] and CMU Sphinx [44]. This design required extensive engineering effort, heavily relied on handcrafted configurations, and lacked the capacity to integrate knowledge holistically or support multitask generalization.
Moreover, in the SLU domain, spoken language understanding was conventionally built on top of ASR outputs. Early SLU systems adopted the ASR $\rightarrow$ NLU pipeline [4, 45], where natural language understanding models operated on 1-best or N-best ASR hypotheses to perform tasks such as intent classification, slot filling, and dialogue state tracking [46]. However, such cascaded systems suffered from error propagation and degraded performance in noisy or ambiguous speech scenarios. Moreover, rich prosodic and acoustic cues were often lost during transcription, limiting deeper semantic interpretation.
Other models for specific speech understanding tasks, such as emotion recognition, have similarly undergone a modular development phase. Although these approaches provide great interpretability, they still exhibit substantial potential for improvement in performance.
Figure 3: The structural evolution of Speech LLMs, illustrating the transition from modular architectures to end-to-end frameworks. Early modular systems is exemplified by traditional ASR pipelines. The end-to-end paradigm subsequently diverges into two branches: E2E Specialized Speech Understanding System and E2E General Speech Understanding System. E2E General Speech Understanding system, also regarded as LLM-Centric systems, are further categorized into discrete sequence modeling (top) and continuous sequence modeling (bottom), reflecting their strategies of representing features. This paper focuses exclusively on understanding-oriented Speech LLMs, as indicated by the blue and blue-gradient blocks.
@@ III-A2 E2E Specialized Speech Understanding System
Complementing these architectural advancements, the E2E era also witnessed significant perceptual progress through self-supervised learning (SSL). Early hand-crafted features such as MFCCs were gradually replaced by rich, contextualized representations learned by models such as wav2vec 2.0 [23], HuBERT [24], and WavLM [25]. These SSL-based encoders substantially enhanced robustness and generalization, and have become foundational in many state-of-the-art E2E systems [49]. However, despite these improvements, conventional E2E models often rely on relatively shallow decoders with limited linguistic capacity, making them less suitable for tasks that require long-range reasoning or rich semantic interpretation.
@@ III-A3 E2E General Speech Understanding System
To address the need for richer and more general speech understanding tasks, a more recent trend within the end-to-end paradigm is the development of E2E General Speech Understanding Systems or LLM-Centric Systems, which integrate LLMs as core reasoning components. Speech understanding models centered around LLMs reflect a paradigm shift toward cognition-guided perception. This trend marks an important development in the evolution of multimodal systems, where the language model no longer holds an equal status with the perceptual front-end, but rather serves as the central controller of interpretation. The conceptual roots of this perspective can be traced back to [51]. These systems represent a structural and functional evolution from conventional E2E models: shifting the focus from acoustic-driven generation to language-model-driven interpretation. Unlike traditional AED-based architectures, LLM-Centric systems align speech inputs to the text modality and utilize LLMs for decoding, generation, and reasoning [26, 30, 27, 31]. This cross-modal design not only enables transfer to diverse downstream tasks—such as summarization, translation, and question answering—but also brings enhanced modularity, interpretability, and instruction-following capability [28, 33, 32, 29].
@@
The evolution from modular pipelines to end-to-end systems, and further into LLM-Centric frameworks, charts a trajectory of increasing unification between wide range of speech understanding tasks like ASR and SLU. While ASR emphasizes accurate linguistic transduction and SLU targets semantic reasoning, both perspectives have progressively converged under the umbrella of Speech LLMs. These models offer a shared representational space and a unified decoding mechanism, making it possible to perform transcription, understanding, and generation in a cohesive framework. In this light, the emergence of LLM-Centric Speech LLMs marks a pivotal point in the development of general-purpose spoken language understanding—a capability referred to as Speech Understanding (SU) in this paper, as illustrated in Section II in detail.
@@
Representative works in this line of research include VALL-E[57], which leverages EnCodec tokens for zero-shot text-to-speech synthesis; SpeechGPT[58], which trains LLMs on unit-text pairs to allow the model to ’listen’ and ’speak’; AudioPaLM[59], which combines wav2vec-style tokenization with multimodal pretraining; VQTTS[60], which introduces self-supervised VQ acoustic features to bridge the gap between ground-truth and predicted features, leading to improved synthesis quality, among other recent methods[61, 62].
This modeling approach offers several advantages. It allows speech processing to benefit directly from LLM pretraining by transforming the problem into a format that LLMs are optimized for: token sequences. Moreover, it enables bi-directional generation—speech-to-text and text-to-speech—within the same unified framework, making it especially suitable for future research on unified speech-language generation. However, as this paper primarily focuses on speech understanding rather than generation, the latter will not be discussed in detail here.
@@
Continuous modeling has proven particularly effective for applications requiring fine-grained acoustic discrimination, including ASR [31], ST [65]. Models like FireRedASR [31] and LLAST [65] show that state-of-the-art performance can be achieved with minimal architectural modification—often a simple projection layer suffices—and limited fine-tuning [30, 27].
In addition to high accuracy and robustness, this paradigm also addresses structural limitations of conventional systems by unifying perception and reasoning within a single framework [66, 67]. Given its scalability and its alignment with the trend of reusing pretrained language models, continuous modeling has become an increasingly important approach in recent Speech LLM research. This paper, therefore, primarily focuses on continuous sequence modeling, while offering comparative insights into discrete token-based alternatives.
@@
WavPrompt [68] Wav2Vec 2.0 Base Linear GPT-2 (124M)
Pengi CLAP Audio Transformer Linear + Gelu + Dropout + Layernorm GPT-2 (124M)
@@
SALMONN Whisper-Large-v2 + BEATs Encoder Q-Former Vicuna-13B
Qwen-Audio Whisper-Large-v2 Encoder Linear Qwen-7B
Qwen2-Audio Whisper-Large-v3 Encoder Avg-Pool + Linear Qwen2-7B
SLAM-LLM Configurable Encoder1 Configurable Projector1 Configurable LLM1
Mala-ASR WavLM-Large-v2 Encoder Linear + ReLU + Linear LLaMA-based
@@
SeedASR LUISE Encoder A Certain Projector3 A Pretrained LLM3
FireRedASR-LLM Conformer Encoder Linear + ReLU + Linear Qwen2-Instruct-7B
E-chat [75] HuBERT Encoder Transformer Baichuan2-7B-Chat
OSUM Whisper-Medium Encoder 1-D CNN + Transformer Qwen2-7B
Kimi-Audio [77] Whisper-Large-v3 EN + GLM4 Tokenizer4 A Certain Downsampler3 7B Pretrained LLM3
LLaST [78] mHuBERT + Whisper-Large Encoder MLP LLaMA2-7B-Instruct
Speech LLaMA [79] Light Audio Encoder5 CTC Compression + Linear Projection LLaMA-7B
LLM-ST [80] Whisper-Large-v2 Encoder6 GPT-3-13B
LegoSLM [81] USM CTC Layer Gemma-2B
@@
2 Details about specific pre-trained models are not publicly disclosed.
@@
6 Whisper-Large-v2 Encoder without a 30-second padding limit.
For models where specific implementation details are not publicly available through the official GitHub repositories or described in the original papers, the corresponding entries are left unspecified in the table.
@@ IV Model structure
To further understand how Speech LLMs process and reason over spoken input, we examine their underlying architectural design. In particular, we investigate the overall structure of Speech LLMs, which typically follows a modular pipeline comprising distinct stages of processing. Various Speech LLMs architectures have been developed up to now, all of which are structured around three fundamental stages: Modality Feature Extraction, Modality Information Fusion, and LLM Inference, as illustrated in Figure 4.
All Speech LLMs begin by encoding the input information, regardless of the modality. Typically, these Speech LLMs leverage pretrained audio and text encoders to obtain the representations of both modalities. The encoded audio and text embeddings are then fused to prepare them for input into the language model. This step transforms the outputs of the feature extraction process into a format compatible with the language model’s input requirements. At present, most Speech LLMs focus on generating text output; therefore, the LLM inference process typically revolves around a text decoder that receives the fused multimodal features as tokenized input to produce text. For models that optionally support audio output, an LLM decoder is trained to generate audio tokens, which are subsequently converted into speech using a vocoder. However, in this survey, we only focus on the Speech LLMs with text output.
@@ IV-A Modality Feature Extraction
As mentioned in the previous section, during the initial feature extraction stage, the audio is processed in two distinct ways. These approaches primarily differ in the output representation format of the audio features. Continuous Sequence Modeling, which is more commonly adopted in current Speech LLMs, encodes audio using a pre-trained and fine-tuned audio encoder, producing an audio embedding $Z_{a}\in\mathbb{R}^{B\times T\times D}$ where $B$, $T$, and $D$ represent batch size, number of frames, and embedding dimension. In contrast, Discrete Sequence Modeling discretizes the input audio signal into a sequence of discrete audio tokens—comparable to text tokens derived from raw text—which are not represented in embedding form. In this section, we summarize common models and techniques used to extract audio features using these two approaches.
@@ IV-A1 Continuous Sequence Modeling
Continuous sequence modeling is generally more straightforward and typically encodes audio in the form of a time-frequency representation extracted using traditional signal processing-based feature extraction methods from the raw audio signal. Given an audio input $X_{a}$, a pre-trained audio encoder is applied to generate the audio features:
$Z_{a}=g(X_{a}),$ (1)
where $g$ denotes the audio encoding process. The most commonly used audio encoders in mainstream Speech LLMs are Whisper [48] and Conformer [63]. Whisper is a sequence-to-sequence Transformer model trained on a wide range of speech processing tasks, and its encoder is recognized as one of the most powerful for audio input processing. Conformer is another widely adopted architecture that combines convolutional layers with Transformers to effectively model both local and global dependencies in speech signals, particularly for automatic speech recognition tasks. Additional encoders used in mainstream Speech LLMs include WavLM [25], a predictive, self-supervised learning (SSL) pre-trained model. In this stage, aside from the audio encoder itself, subsampling modules are often applied to reduce the temporal resolution of the input features, thereby decreasing computational costs and enabling the model to focus on higher-level representations. Ultimately, the output audio feature embeddings $Z_{a}\in\mathbb{R}^{B\times T\times D}$, serving as the input for the next stage. In the context of continuous sequence modeling, Table I provides a detailed summary of representative models, organized according to their architectural components.
@@ IV-A2 Discrete Sequence Modeling
Discrete sequence modeling converts raw audio into a sequence of discrete tokens that represent acoustic content and can typically be decoded back into high-quality audio. The core idea behind generating discrete tokens lies in vector quantization. Building on VQ-VAE [82], which introduced the concept of encoding continuous audio features into symbolic representations via a learned codebook, current mainstream methods for generating discrete tokens include self-supervised pretrained audio tokenizers such as HuBERT [24], and neural codec models such as EnCodec [55, 83]. A general process involves first encoding raw audio into a sequence of latent representations, as in continuous sequence modeling: $Z_{a}=g(X_{a})$, where $Z_{a}=[\textbf{z}_{1},\textbf{z}_{2},\dots,\textbf{z}_{T}]$ with $\textbf{z}_{t}\in\mathbb{R}^{d}$. Next, a quantization function $q(\cdot)$ is applied to obtain the final discrete audio tokens:
@@
Here, $\mathcal{Q}$ denotes the quantization function that maps continuous vectors to codebook vectors $\textbf{e}_{u_{t}}\in\mathbb{R}^{d}$, resulting in $u_{t}$, the discrete token index from the vocabulary set. The final output is a sequence of discrete audio token indices: $U=[u_{1},u_{2},\dots,u_{T}]$. Specifically, self-supervised pretrained audio tokenizers, such as HuBERT, learn contextualized audio representations via masked prediction tasks and discretize intermediate features using k-means clustering. The resulting cluster indices serve as phoneme-like discrete tokens. Neural codec models, such as EnCodec, adopt an encoder–quantizer–decoder architecture in which the audio is compressed into discrete tokens using residual vector quantization and later reconstructed [84]. These models are trained end-to-end and produce high-fidelity audio that is well-suited for generative tasks. The resulting discrete audio tokens are processed alongside tokenized text, with the integration method detailed in the following section.

### v6

@@ III-A1 Modular Architecture
Early efforts in speech processing largely followed a modular architecture. For example, in ASR, the pipeline typically involved a separately trained acoustic model (AM), language model (LM), and decoder, integrated via weighted finite-state transducers (WFSTs). These components were optimized independently and lacked task adaptability. Typical implementations include Gaussian Mixture Model - Hidden Markov Model (GMM-HMM) frameworks and later Deep Neural Network (DNN-HMM) hybrids [6, 45], which formed the foundation of many large-scale ASR toolkits such as Kaldi [46] and CMU Sphinx [47]. This design required extensive engineering effort, heavily relied on handcrafted configurations, and lacked the capacity to integrate knowledge holistically or support multitask generalization.
Moreover, in the SLU domain, spoken language understanding was conventionally built on top of ASR outputs. Early SLU systems adopted the ASR $\rightarrow$ NLU pipeline [4, 48], where natural language understanding models operated on 1-best or N-best ASR hypotheses to perform tasks such as intent classification, slot filling, and dialogue state tracking [49]. However, such cascaded systems suffer from error propagation and degraded performance in noisy or ambiguous speech scenarios. In addition, rich prosodic and acoustic cues were often lost during transcription, limiting deeper semantic interpretation.
Other models for specific speech understanding tasks, such as emotion recognition, have similarly undergone a modular development phase. Although these approaches provide strong interpretability, they still exhibit substantial potential for improvement in performance.
Figure 3: The structural evolution of Speech LLMs, illustrating the transition from modular architectures to end-to-end frameworks. Early modular systems are exemplified by traditional ASR pipelines. The end-to-end paradigm subsequently diverges into two branches: E2E Specialized Speech Understanding System and E2E General Speech Understanding System. E2E General Speech Understanding systems, also regarded as LLM-Centric systems, are further categorized into discrete sequence modeling (top) and continuous sequence modeling (bottom), reflecting their strategies of representing features.
@@ III-A2 E2E Specialized Speech Understanding System
Complementing these architectural advancements, the E2E era also witnessed significant perceptual progress through self-supervised learning (SSL). Early hand-crafted features such as MFCCs were gradually replaced by rich, contextualized representations learned by models such as wav2vec 2.0 [23], HuBERT [24], and WavLM [25]. These SSL-based encoders substantially enhanced robustness and generalization and became foundational in many state-of-the-art E2E systems [52]. However, despite these improvements, conventional E2E models often rely on relatively shallow decoders with limited linguistic capacity, making them less suitable for tasks that require long-range reasoning or rich semantic interpretation.
@@ III-A3 E2E General Speech Understanding System
To address the need for richer and more general speech understanding tasks, a more recent trend within the end-to-end paradigm is the development of E2E General Speech Understanding Systems or LLM-Centric Systems, which integrate LLMs as core reasoning components. Speech understanding models centered around LLMs reflect a paradigm shift toward cognition-guided perception. This trend marks an important development in the evolution of multimodal systems, where the language model no longer holds an equal status with the perceptual front-end but instead serves as the central controller of interpretation. The conceptual roots of this perspective can be traced back to [54]. These systems represent a structural and functional evolution from conventional E2E models, shifting the focus from acoustic-driven generation to language-model-driven interpretation. Unlike traditional AED-based architectures, LLM-Centric systems align speech inputs to the text modality and utilize LLMs for decoding, generation, and reasoning [26, 30, 27, 31]. This cross-modal design not only enables transfer to diverse downstream tasks—such as summarization, translation, and question answering—but also brings enhanced modularity, interpretability, and instruction-following capability [28, 33, 32, 29].
@@
The evolution from modular pipelines to end-to-end systems, and further into LLM-Centric frameworks, charts a trajectory of increasing unification across a wide range of speech understanding tasks like ASR and SLU. While ASR emphasizes accurate linguistic transduction and SLU targets semantic reasoning, both perspectives have progressively converged under the umbrella of Speech LLMs. These models offer a shared representational space and a unified decoding mechanism, making it possible to perform transcription, understanding, and generation in a cohesive framework. In this light, the emergence of LLM-Centric Speech LLMs marks a pivotal point in the development of general-purpose spoken language understanding—a capability referred to as Speech Understanding (SU) in this paper, as illustrated in Section  II in detail.
@@
Representative studies that model speech purely via discretized acoustic tokens include SpeechGPT and AudioGPT [60, 61]. These methods quantize continuous waveforms into discrete code sequences (e.g., with neural codecs) and train a generative language model over the token streams, enabling speech understanding and synthesis within a unified, text-like modeling framework.
This modeling approach offers several advantages. It allows speech processing to benefit directly from LLM pretraining by transforming the problem into a format that LLMs are optimized for: token sequences. Moreover, it enables understanding and generation within the same unified framework. However, models that rely solely on discrete token representations to encode speech information tend to exhibit degraded perceptual quality in speech perception. First, their performance on semantic information perception tasks such as ASR is often inferior to that of traditional models like Whisper or HuBERT. Second, these models also show limited capability in perceiving fine-grained paralinguistic cues and other subtle aspects of speech. Building upon these considerations, this work does not primarily focus on purely discrete sequence modeling approaches.
@@
Continuous modeling has proven particularly effective for applications requiring fine-grained acoustic discrimination, including ASR [31], ST [63]. Models like FireRedASR [31] and LLAST [63] show that state-of-the-art performance can be achieved with minimal architectural modification. In many cases, a simple projection layer is sufficient, together with limited fine-tuning [30, 27].
Although this paradigm can achieve high accuracy in specific tasks, it also introduces certain challenges. When the amount of training data or task diversity is limited, the model is prone to overfitting [64, 65]. Moreover, when integrating with generative modules, it often requires stronger data-driven support or assistance from discrete modeling approaches  [66, 67]. Given its scalability and its alignment with the trend of reusing pretrained language models [68, 69], continuous modeling has become an increasingly important approach in recent Speech LLM research. This paper, therefore, primarily focuses on continuous sequence modeling, while offering comparative insights into discrete token-based alternatives.
@@
Speech LLM with text-only output
WavPrompt [70] Wav2Vec 2.0 Base Encoder Linear GPT-2 (124M)
Pengi [32] CLAP Audio Transformer Linear + Gelu + Dropout + Layernorm GPT-2 (124M)
@@
SALMONN [26] Whisper-Large-v2 + BEATs Encoder Q-Former Vicuna-13B
Speech LLaMA [72] Light Audio Encoder1 CTC Compression + Linear LLaMA-7B
@@
LLM-ST [74] Whisper-Large-v2 Encoder2 GPT-3-13B
Qwen-Audio [37] Whisper-Large-v2 Encoder Linear Qwen-7B
SLAM-ASR [75] Configurable Encoder3 Configurable Projector3 Configurable LLM3
Audio Flamingo [76] ClapCap Encoder Gated Cross Attention OPT-IML-MAX-1.3B
@@
Speechverse [78] A Pretrained Audio Encoder4 1-D CNN A Pretrained LLM4
Mala-ASR [69] WavLM-Large-v2 Encoder Linear + ReLU + Linear LLaMA-based
Qwen2-Audio [38] Whisper-Large-v3 Encoder Avg-Pool + Linear Qwen2-7B
SeedASR [27] LUISE Encoder A Certain Projector3 A Pretrained LLM3
E-chat [79] HuBERT Encoder Transformer Baichuan2-7B-Chat
LLaST [80] mHuBERT + Whisper-Large Encoder MLP LLaMA2-7B-Instruct
@@
FireRedASR-LLM [31] Conformer Encoder Linear + ReLU + Linear Qwen2-Instruct-7B
OSUM [33] Whisper-Medium Encoder 1-D CNN + Transformer Qwen2-7B
Audio Flamingo 2 [83] AF-CLAP Encoder Gated Cross Attention Qwen2.5-3B
LegoSLM [84] USM Encoder CTC Layer Gemma-2B
DeSTA2.5-Audio [85] Whisper-large-v3 Encoder Q-Former Llama-3.1-8B-Instruct
MiDashengLM [86] Dasheng Audio Encoder MLP Qwen2.5-Omni-7B Thinker
Speech LLM with text & speech output
MinMo [66] SenseVoice-large Encoder Transformer + CNN Qwen2.5-7B-instruct
Qwen2.5-Omni [87] Whisper-large-v3–based Encoder LayerNorm + Avg-Pool + Linear Qwen2.5-7B
Kimi-Audio [88] Whisper-Large-v3 EN + GLM4 Tokenizer6 A Certain Downsampler5 7B Pretrained LLM5
Audio Flamingo 3 [89] AF-Whisper Audio Encoder Qwen-2.5-7B
Step-Audio 2 [67] Custom Audio Encoder7 Custom Adapter7 Custom LLM7
MiMo-Audio [90] Custom Audio Encoder8 Custom Adapter8 MiMo-7B-Base
UALM [91] AF-Whisper Audio Encoder MLP Qwen2.5-7B
Qwen3-Omni [92] AuT9 Custom Adapter9 MoE Transformer9
2 Whisper-Large-v2 Encoder without a 30-second padding limit.
@@
4 Details about specific pretrained models are not publicly disclosed.
@@
7 Step-Audio 2 uses a frozen pretrained latent audio encoder with a trainable 2x downsampling adaptor, and a custom multimodal LLM decoder that outputs interleaved text and audio tokens.
8 Uses the MiMo-Audio-Tokenizer’s 32-layer Transformer audio encoder with RVQ-based discretization as the adapter.
9 AuT refers to a custom Audio Transformer. The adapter is not clearly specified, while the decoder is implemented as a custom MoE Transformer.. For models where specific implementation details are not publicly available through the official GitHub repositories or described in the original papers, the corresponding entries are left unspecified in the table.
@@ IV Model structure
To further understand how Speech LLMs process and reason over spoken input, we examine their underlying architectural design. In particular, we investigate the overall structure of Speech LLMs, which typically follows a modular pipeline comprising distinct stages of processing. Various Speech LLM architectures have been developed up to now, all of which are structured around three fundamental stages: Modality Feature Extraction, Modality Information Fusion, and LLM Inference, as illustrated in Figure 4.
All Speech LLMs begin by encoding the input information, regardless of the modality. Typically, these Speech LLMs leverage pretrained audio and text encoders to obtain representations of both modalities. The encoded audio and text embeddings are then fused to prepare them for input into the language model. This step transforms the outputs of the feature extraction process into a format compatible with the language model’s input requirements. For Speech LLMs that generates text output, which are also the primary focus of this survey, the LLM inference process typically revolves around a text decoder that receives the fused multimodal features as tokenized input to produce text. For models that optionally support audio output, an LLM decoder is trained to generate audio tokens, which are subsequently converted into speech using a vocoder. However, in this survey, we only focus on the Speech LLMs with text output.
@@ IV-A Modality Feature Extraction
As mentioned in the previous section, during the initial feature extraction stage, the audio is processed in two distinct ways. These approaches primarily differ in the output representation format of the audio features. Continuous Sequence Modeling, which is more commonly adopted in current Speech LLMs, encodes audio using a pretrained and fine-tuned audio encoder, producing an audio embedding $\mathbf{Z}_{a}\in\mathbb{R}^{B\times T\times D}$ where $B$, $T$, and $D$ represent batch size, number of frames, and embedding dimension. In contrast, Discrete Sequence Modeling discretizes the input audio signal into a sequence of discrete audio tokens—comparable to text tokens derived from raw text—which are not represented in embedding form. In this section, we summarize common models and techniques used to extract audio features using these two approaches.
@@ IV-A1 Continuous Sequence Modeling
Continuous sequence modeling is generally more straightforward and typically encodes audio in the form of a time-frequency representation extracted using traditional signal processing-based feature extraction methods from the raw audio signal. Given an audio input $X_{a}$, a pretrained audio encoder is applied to generate the audio features:
$\mathbf{Z}_{a}=g(\mathbf{X}_{a}),$ (1)
where $g$ denotes the audio encoding process. The most commonly used audio encoders in mainstream Speech LLMs are Whisper [51] and Conformer [62]. Whisper is a sequence-to-sequence Transformer model trained on a wide range of speech processing tasks, and its encoder is recognized as one of the most powerful for audio input processing. Conformer is another widely adopted architecture that combines convolutional layers with Transformers to effectively model both local and global dependencies in speech signals, particularly for automatic speech recognition tasks. Additional encoders used in mainstream Speech LLMs include WavLM [25], a predictive, self-supervised learning (SSL) pretrained model. In this stage, aside from the audio encoder itself, subsampling modules are often applied to reduce the temporal resolution of the input features, thereby decreasing computational costs and enabling the model to focus on higher-level representations. Ultimately, the output audio feature embeddings $\mathbf{Z}_{a}\in\mathbb{R}^{B\times T\times D}$ serve as the input for the next stage. In the context of continuous sequence modeling, Table I provides a detailed summary of representative models, organized according to their architectural components.
@@ IV-A2 Discrete Sequence Modeling
Discrete sequence modeling converts raw audio into a sequence of discrete tokens that represent acoustic content and can typically be decoded back into high-quality audio. The core idea behind generating discrete tokens lies in vector quantization. Building on VQ-VAE [93], which introduced the concept of encoding continuous audio features into symbolic representations via a learned codebook, current mainstream methods for generating discrete tokens include self-supervised pretrained audio tokenizers such as HuBERT [24], and neural codec models such as EnCodec [58, 94]. A general process involves first encoding raw audio into a sequence of latent representations, as in continuous sequence modeling: $\mathbf{Z}_{a}=g(\mathbf{X}_{a})$, where $\mathbf{Z}_{a}=[\textbf{z}_{1},\textbf{z}_{2},\dots,\textbf{z}_{T}]$ with $\textbf{z}_{t}\in\mathbb{R}^{d}$. Next, a quantization function $q(\cdot)$ is applied to obtain the final discrete audio tokens:
@@
Here, $\mathcal{Q}$ denotes the quantization function that maps continuous vectors to codebook vectors $\textbf{e}_{u_{t}}\in\mathbb{R}^{d}$, resulting in $u_{t}$, the discrete token index from the vocabulary set. The final output is a sequence of discrete audio token indices: $U=[u_{1},u_{2},\dots,u_{T}]$. Specifically, self-supervised pretrained audio tokenizers, such as HuBERT, learn contextualized audio representations via masked prediction tasks and discretize intermediate features using k-means clustering. The resulting cluster indices serve as phoneme-like discrete tokens. Neural codec models, such as EnCodec, adopt an encoder–quantizer–decoder architecture in which the audio is compressed into discrete tokens using residual vector quantization and later reconstructed [95]. These models are trained end-to-end and produce high-fidelity audio that is well suited for generative tasks. The resulting discrete audio tokens are processed alongside tokenized text, with the integration method detailed in the following section.